OpenStudy (anonymous):

[(3 + 9) ÷ (3 × 2)] + [(2 × 12) ÷ 4] I came up with 8

hero (hero):

[(3 + 9) ÷ (3 × 2)] + [(2 × 12) ÷ 4] = [(12) ÷ (6)] + [(24) ÷ 4] = [2 + 6] = 8 Correct